# Learning What to Write: Write-Gated KV for Efficient Long-Context Inference

### v1

@@ 6 Related Work
Studies on attention dynamics reveal that token utility is structurally skewed across heads and positions. streamingllm identify “attention sinks,” demonstrating that retaining a few initial tokens allows for stable streaming inference without the full history. Building on this, fastgen propose profiling attention heads to statically categorize them for distinct retention policies. Mechanistic studies (retrievalhead) further reveal that specific “retrieval heads” are uniquely critical for long-range dependencies. Leveraging these characteristics, duoattention, razorattention, and prulong enhance efficiency by statically identifying retrieval-critical heads to retain full context, while restricting others to local context. Unlike these static heuristics, our approach introduces a learnable, input-dependent admission mechanism that captures dynamic utility patterns at runtime.
@@
A large body of work accelerates long-context inference by selectively attending to a sparse, query-dependent subset of cached KV pairs at each step. These methods approximate attention while maintaining the complete underlying cache for future queries. Existing approaches exploit sparsity in various dimensions, including temporal (infllm; hip; specache), channel (sparq), inter-layer (omnikv), and low-rank (sea; loki) structures. Some retrieval-oriented designs treat the KV cache as an indexed memory (retrievalattn; squeezedattn; retroinfer; pqcache; clusterkv; arkvale; magicpig). Systems-oriented designs further integrate selection with KV-aware paging (quest; lserve), runtime management (infinigen; shadowkv), or learnable indexers (deepseekv32) to scale fine-grained selection to very long contexts. Our write-time admission mechanism is complementary to these read-time strategies, allowing for compound efficiency gains by filtering noise prior to selection.
@@
Another line of work focuses on maintaining a bounded memory by retrospectively pruning previously written states. Representative methods employ statistics derived from past attention or recency to evict low-utility tokens (h2o; tova; scissorhand; keyformer; sirllm; l2norm). Other approaches refine the eviction criteria through learnable interaction scores (dynamiccontextpruning; attentiongate), importance estimation (trimkv; locret), reconstruction objectives (kvzip), or prompt-guided heuristics (snapkv; nacl; pyramidinfer; finch). Unlike admission, which filters intrinsic noise at the source, eviction manages cache bounds by removing obsolete history; thus, the two serve distinct, complementary roles in the token lifecycle.

### v2

@@ 6 Related Work
Studies on attention dynamics reveal that token utility is structurally skewed across heads and positions. streamingllm identify “attention sinks,” demonstrating that retaining a few initial tokens allows for stable streaming inference without the full history. Building on this, fastgen propose profiling attention heads to statically categorize them for distinct retention policies. Mechanistic studies (retrievalhead) further reveal that specific “retrieval heads” are uniquely critical for long-range dependencies. Leveraging these characteristics, razorattn, duoattn, and prulong enhance efficiency by statically identifying retrieval-critical heads to retain full context, while restricting others to local context. Unlike these static heuristics, our approach introduces a learnable, input-dependent admission mechanism that captures dynamic utility patterns at runtime.
@@
A large body of work accelerates long-context inference by selectively attending to a sparse, query-dependent subset of cached KV pairs at each step. These methods approximate attention while maintaining the complete underlying cache for future queries. Existing approaches exploit sparsity in various dimensions, including temporal (infllm; hip; specache), channel (sparq), inter-layer (omnikv), and low-rank (sea; loki) structures. Some retrieval-oriented designs treat the KV cache as an indexed memory (arkvale; retrievalattn; squeezedattn; retroinfer; pqcache; clusterkv; magicpig). Systems-oriented designs further integrate selection with KV-aware paging (quest; lserve), runtime management (infinigen; shadowkv), sparse prefill patterns (minference; flexprefill; xattention), or learnable indexers (nsa; deepseekv32) to scale fine-grained selection to very long contexts. Our write-time admission mechanism is complementary to these read-time strategies, allowing for compound efficiency gains by filtering noise prior to selection.
@@
Another line of work focuses on maintaining a bounded memory by retrospectively pruning previously written states. Representative methods employ statistics derived from past attention or recency to evict low-utility tokens (h2o; scissorhand; tova; keyformer; sirllm; l2norm). Other approaches refine the eviction criteria through learnable interaction scores (dynamiccontextpruning; attentiongate), importance estimation (trimkv; locret; dms), reconstruction objectives (kvzip), or prompt-guided heuristics (snapkv; nacl; pyramidinfer; finch). However, unlike our hardware-aware admission mechanism, these methods lack specialized memory management for the resulting ragged KV states and do not address the quadratic complexity of the prefill phase, limiting their realizable end-to-end speedup.